A roof rises 2.75 ft over a horizontal distance of 15.01 ft. What is the slope of the roof to the nearest hundredth?
Thank you.

Answer by checkley71(8403)

USING A REFERENCE POINT OF (0,0) AND THE EDGE OF THE ROOF IS (15.01,2.75)
WE HAVE THE FOLLOWING SLOPE:
(2.75-0)/(15.01-0)=2.75/15.01=.183 IS THE SLOPE.